An investor currently holds the following portfolio: Amount Invested 8,000 shares of Stock A $16,000 Beta = 1.3 15,000 shares of Stock B $48,000 Beta = 1.8 25,000 shares of Stock C $96,000 Beta = 2.2 If the risk-free rate of return is 2% and the market risk premium is 7%, then the required return on the portfolio is A. 14.91%. B. 23.93%. C. 21.91%. D. 15.93%.
Answer #1

Beta of portfolio = 1.3*16000/(16000+48000+96000) +1.8*48000/(16000+48000+96000) + 2.2*96000/(16000+48000+96000)

Beta of portfolio = 1.99

Required return on the portfolio = Rf + market risk premium * Beta

Required return on the portfolio = 2 + 7*1.99

Required return on the portfolio = 15.93%

Answer

D. 15.93%.
Answer #2

Total amount: 16000+48000+96000 = 160,000,

Weight of A: 16000/160000=10%;

Weight of B: 48000/160000=30%;

Weight of A: 96000/1600000=60%;

So weighted average of beta is:

Weight of A * beta of A + Weight of B * beta of B + Weight of C * beta of C

= 10% * 1.3 + 30% * 1.8 + 60% * 2.2 = 1.99

then the required return on the portfolio is

A. 14.91%.

B. 23.93%.

C. 21.91%

. D. 15.93%.--answer